OpenStudy (anonymous):

The x-coordinates of two objects moving along the x-axis are given as a function of time (t). x1= (4m/s)t x2= -(161m) + (48m/s)t - (4 m/s^2)t^2 Calculate the magnitude of the distance of closest approach of the two objects. x1 and x2 never have the same value.

OpenStudy (asnaseer):

create an equation to represent the distance between the two x's. you can do this by finding x1-x2. lets call this 'y':\[y=x_1-x_2=4t-(-161+48t-4t^2)=4t+161-48t+4t^2\]\[y=4t^2-44t+161\]we want to minimize the value of y. this will happen when the first derivative of y is zero. so:\[y'=8t-44=0\implies t=\frac{44}{8}=\frac{11}{4}\]substitute this value for t into the equation for y to get the distance between them as follows:\[y_{min}=4(\frac{11}{4})^2-44(\frac{11}{4})+161=\frac{121}{4}-121+161=9.75m\]

OpenStudy (anonymous):

This is not the correct answer....

OpenStudy (asnaseer):

what do you think is the correct answer?

OpenStudy (anonymous):

Well I'm given an array of multiple choice answers...

OpenStudy (asnaseer):

which are?

OpenStudy (anonymous):

1. 39 m 2. 57 m 3. 43 m 4. 29 m 5. 35 m 6. 61 m 7. 27 m 8. 46 m 9. 34 m 10. 40 m

OpenStudy (asnaseer):

are you certain that you have stated the question correctly?

OpenStudy (anonymous):

The x-coordinates of two objects moving along the x-axis are given as a function of time t. x1 = (4 m/s) t and x2 = −(161 m) + (48 m/s) t − (4 m/s2) t2 . Calculate the magnitude of the distance of closest approach of the two objects. x1 and x2 never have the same value.

OpenStudy (asnaseer):

then I cannot see where the mistake is - can you?

OpenStudy (anonymous):

The calculation doesnt seem to appear right in the last part. It equas over a hundred... that isnt correct either.

OpenStudy (anonymous):

equals*

OpenStudy (anonymous):

oh 11/2 is t

OpenStudy (asnaseer):

Oops - I added up incorrectly at the end, it should have been:\[y_{min}=4(\frac{11}{4})^2-44(\frac{11}{4})+161=\frac{121}{4}-121+161=70.25m\]but that doesn't match any of your answers either?

OpenStudy (asnaseer):

aha - yes - well spotted, so we get:\[y_{min}=4(\frac{11}{2})^2-44(\frac{11}{2})+161=\frac{121}{4}-121+161=40m\]

OpenStudy (anonymous):

yay.... crossing my fingers.

OpenStudy (asnaseer):

sorry about my silly mistakes :-(

OpenStudy (anonymous):

I should have actually looked and figured that out myself. But thank you. Do you think you could help me on this hot air balloon problem?

OpenStudy (asnaseer):

have you posted it on the left?

OpenStudy (anonymous):

yes.

OpenStudy (asnaseer):

ok - let me take a look